2x + 5y= 10
50x+ 125y= 100

Answer by Boreal(15235)

2x + 5y= 10
50x+ 125y= 100
multiply top by -25
-50x-125y=-2500, add
0=-2400. no solution. Lines are parallel.
Can write as y=mx+b
5y=-2x+10 first one
y=-(2/5)x+2, dividing both sides by 5
125y=-50x+100
y=(-50/125)x+(100/125) after dividing both sides by 125
y=(-2/5)x+(4/5)
both have the same slope but different y-intercepts. That is the definition of parallel lines.